OpenStudy (anonymous):

How do I graph x^2y - y^3 - 5x^2 + 5y^2 = 0 ? I factored it to get (y-5)(x-y)(x+y) = 0 but would it be correct to graph each factor individually?

OpenStudy (anonymous):

Good one:-) y= plus minus x or 5, lol

OpenStudy (anonymous):

Did u try just doing it as a table? U know, set y at 0, -1,1, etc?

OpenStudy (anonymous):

plugging values for y seems to create the same graph as if i had graphed the factors individually, except that it doesn't extend over to quadrants II and III... if i did that correctly...

OpenStudy (anonymous):

Wolframs version.. http://www.wolframalpha.com/input/?i=x^2y+-+y^3+-+5x^2+%2B+5y^2+%3D+0

OpenStudy (anonymous):

oh ok so would graphing the factors individually work for other equations as well? or are there exceptions?

OpenStudy (anonymous):

I think someone was having a little joke with u when they gave u this one..:-) If u just think of a quadratic equation it is clear that graphing the factors is not correct. That doesn't mean that the factors don't provide useful info, though.

OpenStudy (anonymous):

haha ok thanks :)

OpenStudy (anonymous):

ur welcome.

OpenStudy (anonymous):

I came across this question in a Pre-calculus book. The answer was actually very simple. \[x ^{2}y-y ^{3}-5x ^{2}+5y ^{2}=0\] Then move your isolate your variables like this: \[x ^{2}y-5x ^{2}=y ^{3}-5y^{2}\] \[x ^{2}(y-5)=y ^{2}(y-5)\] Finally, simplify the equation to get: \[x ^{2}=y ^{2}\] x=y

OpenStudy (anonymous):

Cubic in y means 3 roots (not 1) for y, y= plus or minus x and y= 5